Solve 40x is congruent to 777 (mod 1777).
Answer by ikleyn(53763)

.
Answer.  x = 1752.



Check.  40*1752 - 777 = 69303 = 39*1777.



How did I find it? - Using MS Excel in my computer.

                     and also knowing that 40x = 777 (mod 1777) means that the difference 40x - 777 is a multiple of 1777.

Solved.

============================
Dear "students" !

Today is a special day.
Today I solved  16  problems for you and got back  0  (zero,  ZERO)  "thanks"  from you.

It is the absolute record !
    It is just out of limits of my understanding  !
        I never saw it before  !

Thank you very much and wish you to think:   WHO ARE YOU . . .